OpenStudy (anonymous):

solve for x. show all steps leading to the solution 4x+1/3=x-5/2

OpenStudy (anonymous):

Cross multiply 2(4x +1) = 3(x-5) 8x +2 = 3x -15 5x = -17 x = -17/5

OpenStudy (aroub):

x=-17/18

OpenStudy (anonymous):

it would be -3.4 in decimals right?